OpenStudy (anonymous):

A sample of oxygen gas (O2) has a temperature of 273 K at 1.00 atm. What should the density of this gas be

OpenStudy (astrophysics):

\[\rho = \frac{ mass }{ volume }\] where \[\rho = density \] \[P=1 atm = 101 kPa\]\ \[T = 273K\]

OpenStudy (astrophysics):

So we can find the molar mass of oxygen \[M_{O_2} = 32 \frac{ g }{ mol }\] are you given more information?

OpenStudy (unklerhaukus):

Use \[PV=nRT\]and \[n = m/M\]

OpenStudy (jfraser):

as @unklerhaukus says, use the ideal gas law and rearrange it to solve for \(\frac{mass}{volume}\)